OpenStudy (anonymous):

If 100 checks cost $12.00, find the cost per check.

jhonyy9 (jhonyy9):

do you know the rule of 3 simple ?

jhonyy9 (jhonyy9):

if 100 cost 12 so than 1 cost x -------------------- from what x=(1*12)/100 =12/100 =0,12

OpenStudy (anonymous):

Thank you. :) so the answer is .12?

jhonyy9 (jhonyy9):

yes right sure good luck bye

OpenStudy (anonymous):

Thank you. :) God bless you.

jhonyy9 (jhonyy9):

you too bye